### v2

# On homomorphic images of ultraproducts

Abstract.
In this short note we answer some questions of Bergman regarding homomorphic images of (ultra)products of groups.
Key words and phrases:
ultraproduct, algebraically compact group, cotorsion group
2020 Mathematics Subject Classification:
Primary 03C20, 20K25; Secondary 08B25
The work of the author was supported by Basque Government Grant IT1483-22 and Spanish Government Grants PID2019-107444GA-I00 and PID2020-117281GB-I00.
1. Introduction
Ultrafilter constructions provide a rich source of examples and simplify many arguments. Though such constructions nicely serve their raison d’être, they tend to be mysterious and unwieldy in some ways. We consider natural questions in group theory, which derive from the following question. If $\mathcal{U}$ is a nonprincipal ultrafilter over the set $\omega$ of natural numbers, then what can be said regarding homomorphic images of an ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{U}$?
We focus on some specific questions of George Bergman. For Questions 1 and 2 see [1, Qu. 35], [12, Qu. 20.13 (a)] and respectively [12, Qu. 20.13 (b)]. Questions 3 and 4 are respectively [1, Qu. 19] and [1, Qu. 33]. Question 5 is [1, Qu 17], [12, Qu. 20.10 (b)].
Question 1.
If an abelian group can be written as a homomorphic image of a non-principal countable ultraproduct of not necessarily abelian groups $G_{i}$, must it be a homomorphic image of a nonprincipal countable ultraproduct of abelian groups?
Question 2.
If an abelian group can be written as a homomorphic image of a direct product of an infinite family of not necessarily abelian finite groups, can it be written as a homomorphic image of a direct product of finite abelian groups?
Question 3.
If $\mathcal{U}$ is a nonprincipal ultrafilter on $\omega$ and $H$ is a group such that every $h\in H$ lies in a homomorphic image within $H$ of $(\prod_{\omega}\mathbb{Z})/\mathcal{U}$, must $H$ be a homomorphic image of a nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{V}$?
Question 4.
If an abelian group $A$ is such that every homomorphism $\bigoplus_{\omega}\mathbb{Z}\rightarrow A$ extends to a homomorphism $\prod_{\omega}\mathbb{Z}\rightarrow A$, then is $A$ cotorsion (see Definition 6)?
Question 5.
If $\mathcal{U}$ and $\mathcal{U}^{\prime}$ are nonprincipal ultrafilters on $\omega$, does there exist a nonprincipal ultrafilter $\mathcal{U}^{\prime\prime}$ such that every group which is the homomorphic image of an ultraproduct over $\mathcal{U}$ or over $\mathcal{U}^{\prime}$ can be written as an image of an ultraproduct over $\mathcal{U}^{\prime\prime}$?
Questions 1 and 2 will be answered affirmatively using some new characterizations of cotorsion abelian groups (Theorem 17). We’ll see that the direct sum $\bigoplus_{n\in\omega\setminus\{0\}}\mathbb{Z}/n\mathbb{Z}$ gives a negative answer to Question 3, and many more examples will follow from Corollary 23. It was added in the proof of the article [1] that Question 4 has a positive solution; we provide a very short argument using the machinery presented. It may not be immediately clear what Questions 2 and 4 have to do with ultraproducts, but it will be seen that all these questions are of a kind. The principal tool in attacking Questions 1 - 4 is the notion of Higman-completeness, due to Herefort and Hojka [10]. We will not completely answer Question 5, but will show that an affirmative answer is consistent with ZFC. Indeed, a positive answer follows from Blass’s near coherence of filters [2].
The paper is organized as follows. Questions 1 - 4 are answered in Section 2. A greater variety of examples for Question 3 is given in Section 3, and Question 5 is treated in Section 4.
2. Higman-complete groups
We begin this section by defining a handful of concepts and state a proposition detailing some known relationships among them. Next, we pass through some lemmas towards Theorem 17, answering some questions along the way. We emphasize that although much of the discourse will involve abelian groups, we do not require any of our groups to be abelian unless so stated.
Definition 6.
[8, §9.6] An abelian group $A$ is cotorsion if every short exact sequence of abelian groups
$0\rightarrow A\rightarrow B\rightarrow C\rightarrow 0$
with $C$ torsion-free necessarily splits.
Definition 7.
[8, §6.1] An abelian group $A$ is algebraically compact if there exists a compact abelian group $\mathbb{A}$ in which $A$ is a direct summand.
Definition 8.
[10] A group $G$ is Higman-complete if for each sequence $(f_{i})_{i\in\omega}$ of elements in $G$ and sequence of free words in two variables $(w_{i})_{i\in\omega}$ there exists a sequence $(h_{i})_{i\in\omega}$ of elements in $G$ such that $h_{i}=w_{i}(f_{i},h_{i+1})$ for each $i\in\omega$.
Remark 9.
[10, Lemma 2] The class of Higman-complete groups is closed under taking homomorphic images.
The following is known.
Proposition 10.
Let $A$ be an abelian group. The following are equivalent.
  1. (i)
    $A$ is cotorsion.
  2. (ii)
    $A$ is a homomorphic image of an algebraically compact group.
  3. (iii)
    $A$ is Higman-complete.
  4. (iv)
    $A$ is the homomorphic image of an ultraproduct $(\prod_{i\in\omega}A_{i})/\mathcal{U}$, where $\mathcal{U}$ is nonprincipal and the $A_{i}$ are abelian.
Proof.
The equivalence of (i) and (ii) is classically known. The equivalence of (i) and (iii) is [10, Theorem 3]. The equivalence of (i) and (iv) is in [1, Proposition 23]. ∎
With this information we are already equipped for the following.
Proof of affirmative solution to Question 4.
Assume the hypotheses. We will show that $A$ is Higman-complete, and the fact that $A$ is cotorsion then follows immediately from Proposition 10. Let $(f_{i})_{i\in\omega}$ be a sequence of elements of $A$ and $(w_{i})_{i\in\omega}$ a sequence of words in two variables. Since $A$ is abelian, we can without loss of generality assume that the word $w_{i}(x,y)$ is of form $z_{i}x+z_{i}^{\prime}y$ where $z_{i},z_{i}^{\prime}\in\mathbb{Z}$. For each $i\in\omega$, let $\delta_{i}\in\mathbb{Z}^{\omega}$ be the element with $\delta_{i}(i)=1$ and $\delta_{i}(j)=0$ for $j\neq i$. Letting $\phi:\bigoplus_{\omega}\mathbb{Z}\rightarrow A$ be defined by the function $\delta_{i}\mapsto f_{i}$, we obtain (by the hypothesis of Question 4) an extension $\Phi:\mathbb{Z}^{\omega}\rightarrow A$. Letting
  • $h_{0}=\Phi(z_{0},z_{0}^{\prime}z_{1},z_{0}^{\prime}z_{1}^{\prime}z_{2},z_{0}^{\prime}z_{1}^{\prime}z_{2}^{\prime}z_{3},\ldots)$
  • $h_{1}=\Phi(0,z_{1},z_{1}^{\prime}z_{2},z_{1}^{\prime}z_{2}^{\prime}z_{3},\ldots)$
  • $h_{2}=\Phi(0,0,z_{2},z_{2}^{\prime}z_{3},z_{2}^{\prime}z_{3}^{\prime}z_{4},\ldots)$
    $\vdots$
it is easy to check that the sequence $(h_{i})_{i\in\omega}$ satisfies $h_{i}=w_{i}(f_{i},h_{i+1})$. For example,
$h_{0}=\Phi(z_{0}\delta_{0})+z_{0}^{\prime}\Phi(0,z_{1},z_{1}^{\prime}z_{2},\ldots)=z_{0}f_{0}+z_{0}^{\prime}h_{1}$
and for $i>0$ the check is entirely analogous. Thus $A$ is Higman-complete, and therefore cotorsion by Proposition 10.
Lemma 11.
For an arbitrary sequence of groups $(G_{i})_{i\in\omega}$ the group
$G=(\prod_{i\in\omega}G_{i})/(\bigoplus_{i\in\omega}G_{i})$
is Higman complete. Also, for any nonprincipal ultrafilter $\mathcal{U}$ the ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{U}$ is Higman-complete.
Proof.
In this lemma we will use symbols $\zeta$ and $\eta$ to denote elements of $\prod_{i\in\omega}G_{i}$ and $[\zeta],[\eta]$ to denote elements of $G$. Let a sequence $([\zeta_{\ell}])_{\ell\in\omega}$ of elements of $G$ be given, together with a sequence $(w_{\ell})_{\ell\in\omega}$ of free words in two variables. For each $\ell\in\omega$ we fix a representative $\zeta_{\ell}\in[\zeta_{\ell}]$. We shall define by induction a sequence $(\eta_{\ell})_{\ell\in\omega}$ of elements in $\prod_{i\in\omega}G_{i}$. Let $\eta_{0}(0)=w_{0}(\zeta_{0}(0),1_{G_{0}})$. Let $\eta_{1}(1)=w_{1}(\zeta_{1}(1),1_{G_{1}})$ and $\eta_{0}(1)=w_{0}(\zeta_{0}(1),\eta_{1}(1))$. Generally, if $\eta_{\ell}(j)$ has been defined for all $\ell\leq j\leq k$, then we let
  • $\eta_{k+1}(k+1)=w_{k+1}(\zeta_{k+1}(k+1),1_{G_{k+1}})$
  • $\eta_{k}(k+1)=w_{k}(\zeta_{k}(k+1),\eta_{k+1}(k+1))$
  • $\eta_{k-1}(k+1)=w_{k-1}(\zeta_{k-1}(k+1),\eta_{k}(k+1))$
    $\vdots$
  • $\eta_{0}(k+1)=w_{0}(\zeta_{0}(k+1),\eta_{1}(k+1))$.
For $j<\ell$ we let $\eta_{\ell}(j)=1_{G_{j}}$.
We claim that the sequence $([\eta_{\ell}])_{\ell\in\omega}$ of elements in $G$ satisfies the definition of Higman-completeness. To see this, let $\ell\in\omega$ be given. By construction we have for all $j\geq\ell$ that $\eta_{\ell}(j)=w_{\ell}(\zeta_{\ell}(j),\eta_{\ell+1}(j))$, so that we can indeed write $[\eta_{\ell}]=w_{\ell}([\zeta_{\ell}],[\eta_{\ell+1}])$.
For the claim in the second sentence of this lemma we note that $(\prod_{i\in\omega}G_{i})/\mathcal{U}$ is a quotient of $(\prod_{i\in\omega}G_{i})/(\bigoplus_{i\in\omega}G_{i})$ and apply Remark 9.
Proof of negative answer to Question 3.
Consider the abelian group
$H=\bigoplus_{n\in\omega}C_{n}$
where the $C_{n}$ are finite cyclic groups of unbounded orders. This group is torsion, and so each $h\in H$ is an element of a finite cyclic subgroup $C$ of $H$, say $|C|=m$, and it a trivial matter to surject $(\prod_{\omega}\mathbb{Z})/\mathcal{U}$ onto $(\prod_{\omega}\mathbb{Z}/m\mathbb{Z})/\mathcal{U}$ and as $m$ is finite we have $(\prod_{\omega}\mathbb{Z}/m\mathbb{Z})/\mathcal{U}$ isomorphic to $C$. Suppose by way of contradiction that $H$ is a homomorphic image of a nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{V}$. Then $H$ is Higman-complete (by Lemma 11 and Remark 9). Thus $H$ is an abelian group which is Higman-complete, and so $H$ is cotorsion (Proposition 10). Since $H$ is a torsion cotorsion group, it is a direct sum of a bounded group (i.e. a group in which $a^{n}=0$ for some universal $n\in\omega$) and a divisible group ([8, Corollary 9.8.4]). However $H$ has no nontrivial divisible elements and is not bounded, a contradiction. We shall later see that many nonabelian groups of this flavor also serve as counterexamples. ∎
We interrupt at this stage to give a result involving elementary equivalence, which may be of independent interest.
Proposition 12.
Let $G$ be a group. The following are equivalent.
  1. (a)
    There exists a group $H$ which is elementarily equivalent to $G$ such that the abelianization $H/H^{\prime}$ is nontrivial cotorsion.
  2. (b)
    Either $G$ is not perfect or $G$ has elements of unbounded commutator length.
Proof.
(b) $\Rightarrow$ (a). Take $\mathcal{U}$ to be a nonprincipal ultrafilter on $\omega$. Let $H$ be the ultrapower $(\prod_{\omega}G)/\mathcal{U}$. Then $H$ is Higman-complete (Lemma 11), elementarily equivalent to $G$, and the abelianization $H/H^{\prime}$ is cotorsion. It remains to prove that $H/H^{\prime}$ is nontrivial. If $G$ is not perfect, then the natural surjective map from $H=(\prod_{\omega}G)/\mathcal{U}$ to $(\prod_{\omega}G/G^{\prime})/\mathcal{U}$ gives a nontrivial abelian image of $H$, so $H/H^{\prime}$ is nontrivial. On the other hand if $G$ is perfect then $G$ has elements of arbitrarily long commutator length, so select an element $(g_{i})_{i\in\omega}$ in $\prod_{i\in\omega}G$ such that the commutator length of $g_{i}$ is greater than $i$. Then the element $[(g_{i})_{i\in\omega}]$ in $H$ is not in the kernel of the abelianization of $H$ (since its commutator length is infinite).
(a) $\Rightarrow$ (b) We prove the contrapositive. Supposing that $G$ is perfect and of bounded commutator length, say length $j$, every group $H$ which is elementarily equivalent to $G$ also is perfect of bounded commutator length at most $j$ (since this property can be expressed using a first-order formula). Then the abelianization of such an $H$ is trivial. ∎
Continuing our progress toward Theorem 17 we give two more lemmas.
Lemma 13.
If $\{G_{j}\}_{j\in J}$ is a collection of Higman-complete groups then the product
$G=\prod_{j\in J}G_{j}$
is Higman-complete.
Proof.
Given a sequence $(\zeta_{i})_{i}$ of elements in $G$, and a sequence $(w_{i})_{i\in\omega}$ of free words in two variables, we solve the necessary equations coordinatewise. More precisely, for a fixed $j\in J$ we have a sequence $(\zeta_{i}(j))_{i\in\omega}$ of elements of $G_{j}$ and we select $(\eta_{i}(j))_{i\in\omega}$ so that $\eta_{i}(j)=w_{i}(\zeta_{i}(j),\eta_{i+1}(j))$ for each $i\in\omega$ (which is possible since $G_{j}$ is Higman-complete). Now, it is clear that $\eta_{i}=w_{i}(\zeta_{i},\eta_{i+1})$ holds, since the equation holds in each coordinate. ∎
Definition 14.
[8, §4.1] An abelian group is reduced if it has no nontrivial divisible elements.
Lemma 15 below is immediate from [8, Theorem 4.3.1], and Lemma 16 is probably known.
Lemma 15.
An abelian group $D$ is divisible if and only if $D$ is the homomorphic image of $\bigoplus_{\kappa}\mathbb{Q}$ for some cardinal $\kappa$.
Lemma 16.
If $\mathbb{A}$ is an algebraically compact group, then $\mathbb{A}$ is a homomorphic image of a product of finite cyclic groups.
Proof.
Letting $D$ denote the maximal divisible subgroup of $\mathbb{A}$, we can write $\mathbb{A}=D\oplus R$ where $R$ is a reduced algebraically compact group. A reduced algebraically compact group is a direct summand of a product of cyclic groups of prime power order [8, Corollary 6.1.4], so in particular $R$ is a homomorphic image of a product of finite cyclic groups.
Letting $\mathbb{P}$ denote the set of prime natural numbers, it is easy to see that $B=(\prod_{p\in\mathbb{P}}\mathbb{Z}/p\mathbb{Z})/(\bigoplus_{p\in\mathbb{P}}\mathbb{Z}/p\mathbb{Z})$ is torsion-free, divisible, and infinite. Thus $B$ is a vector space over the field $\mathbb{Q}$ of dimension greater than 00, so by picking a basis and projecting to a single coordinate we obtain a homomorphism from $B$ onto $\mathbb{Q}$. In particular, $\mathbb{Q}$ is a homomorphic image of a product of finite cyclic groups.
By Lemma 15 take $\kappa$ to be a cardinal such that $D$ is a homomorphic image of $\bigoplus_{\kappa}\mathbb{Q}$. Now $\bigoplus_{\kappa}\mathbb{Q}$ is a homomorphic image of $\mathbb{Q}^{\kappa}$ (by a vector space argument), and $\mathbb{Q}^{\kappa}$ is a homomorphic image of a product of finite cyclic groups (since $\mathbb{Q}$ is such an image). Thus $D$ is a homomorphic image of a product of finite cyclic groups, and as $R$ also satisfies this property, the group $\mathbb{A}=D\oplus R$ also has this property.
Theorem 17.
Let $A$ be an abelian group. The following are equivalent.
  1. (1)
    $A$ is cotorsion.
  2. (2)
    $A$ is a homomorphic image of $(\prod_{i\in\omega}G_{i})/(\bigoplus_{i\in\omega}G_{i})$ for some sequence $(G_{i})_{i\in\omega}$ of groups which are not necessarily abelian.
  3. (3)
    $A$ is a homomorphic image of a product of finite cyclic groups.
  4. (4)
    $A$ is a homomorphic image of a product $\prod_{j\in J}F_{j}$ where each $F_{j}$ is a finite group which is not necessarily abelian.
Proof.
We will first prove the equivalence of (1) and (2). Suppose that $A$ is cotorsion. By Proposition 10 we know that there is a sequence $(A_{i})_{i\in\omega}$ of abelian groups and a nonprincipal ultrafilter $\mathcal{U}$ on $\omega$ so that $A$ is a homomorphic image of the ultraproduct $(\prod_{i\in\omega}A_{i})/\mathcal{U}$. As $(\prod_{i\in\omega}A_{i})/\mathcal{U}$ is a homomorphic image of $(\prod_{i\in\omega}A_{i})/(\bigoplus_{i\in\omega}A_{i})$, so is $A$. For the other direction, if $A$ is a homomorphic image of a group of form $(\prod_{i\in\omega}G_{i})/(\bigoplus_{i\in\omega}G_{i})$ then $A$ is Higman-complete (Lemma 11), so $A$ is cotorsion (Proposition 10).
Now we argue that (1) $\Rightarrow$ (3) $\Rightarrow$ (4) $\Rightarrow$ (1). That (1) implies (3) follows from the fact that a cotorsion group is a homomorphic image of an algebraically compact group (Proposition 10) and the fact that an algebraically compact abelian group is a homomorphic image of a product of finite cyclic groups (Lemma 16). That (3) implies (4) is evident. For the last implication, note that a finite group $F$ is Higman-complete. To see this, take a nonprincipal ultrafilter $\mathcal{V}$ on $\omega$ and recall that $F$ is isomorphic to the ultrapower $(\prod_{\omega}F)/\mathcal{V}$, so $F$ is Higman-complete (Proposition 11). Therefore a product of finite groups is Higman-complete (Lemma 13), and so $A$ is Higman-complete as the homomorphic image of a Higman-complete group (Remark 9). Thus $A$ is cotorsion (Proposition 10).
Now it is clear that Question 1 has an affirmative solution (using (2) $\Rightarrow$ (1) of Theorem 17 and (i) $\Rightarrow$ (iv) of Proposition 10) as does Question 2 (using (4) $\Rightarrow$ (3)).
3. An analogue of a theorem of Chase
In the light of the very abelian nature of many of the previous results, we give a general method for producing (nonabelian) examples which give a negative solution to Question 3. These will be torsion groups which are not the homomorphic images of any nonprincipal ultraproduct $(\prod_{i\in\omega}G_{i})/\mathcal{U}$. For this we will prove an analogue of a result of Chase [7].
Definitions 18.
We’ll say a group $H$ is a bounded torsion group if there exists some $n\in\omega$ such that $(\forall h\in H)h^{n}=1_{H}$. For such an $H$ we let $\operatorname{ex}(H)\in\omega$ denote the smallest such $n$. If a torsion group is not bounded then we shall say it is an unbounded torsion group.
Definitions 19.
Suppose $\{H_{j}\}_{j\in J}$ is a collection of bounded torsion groups. For $h=(h_{j})_{j\in J}\in\prod_{j\in J}H_{j}$ we let $\textbf{supp}(h)=\{j\in J\mid h_{j}\neq 1_{H_{j}}\}$. Define
$\textbf{Bdd}\{H_{j}\}_{j\in J}=\{h\in\prod_{j\in J}H_{j}\mid(\exists n\in\omega)(\forall j\in\textbf{supp}(h))\operatorname{ex}(H_{j})\leq n\}.$
Remark 20.
It is clear that $\textbf{Bdd}\{H_{j}\}_{j\in J}$ is a subgroup of $\prod_{j\in J}H_{j}$, since if $n$ and $n^{\prime}$ respectively witness that $h$ and $h^{\prime}$ are in $\textbf{Bdd}\{H_{j}\}_{j\in J}$ then $\max(n,n^{\prime})$ witnesses that $hh^{\prime}\in\textbf{Bdd}\{H_{j}\}_{j\in J}$. Moreover, this subgroup is torsion.
Theorem 21.
Let $G$ be a topological group which is either
  1. (1)
    completely metrizable; or
  2. (2)
    Hausdorff and locally countably compact
and each $H_{j}$ a bounded torsion group and let $\phi:G\rightarrow\textbf{Bdd}\{H_{j}\}_{j\in J}$ be an abstract group homomorphism. Then there is some open neighborhood $U$ of the identity $1_{G}$ and $N\in\omega\setminus\{0\}$ such that $(\forall h\in\phi(U))h^{N}=1$.
Proof.
Let $H=\textbf{Bdd}\{H_{j}\}_{j\in J}$. Suppose by way of contradiction that the conclusion fails. The strategy will be to produce a system of equations (like in the definition of Higman-completeness) whose solution will give a contradiction. Suppose we are in situation (1) and let $d$ be a complete metric which induces the topology on $G$. We inductively choose
  1. (a)
    a sequence $(g_{m})_{m\in\omega}$ of elements in $G$; and
  2. (b)
    a sequence $(r_{m})_{m\in\omega}$ of positive natural numbers.
Select $g_{0}\in G$ such that $\phi(g_{0})$ has order greater than $1=0!$. Let $r_{0}$ be the least common multiple of the finite nonempty set $\{\operatorname{ex}(H_{j})|j\in\textbf{supp}(\phi(g_{0}))\}$. Assume we have chosen $g_{0},\ldots,g_{m}$ and $r_{0},\ldots,r_{m}$. Select neighborhood $U$ of $1_{G}$ so that $g\in U$ implies
  • $d(g_{m}g^{r_{m}},g_{m})<\frac{1}{3}$;
  • $d(g_{m-1}(g_{m}g^{r_{m}})^{r_{m-1}},g_{m-1}(g_{m})^{r_{m-1}})<\frac{1}{3^{2}}$;
    $\vdots$
  • $d(g_{0}(g_{1}(\cdots(g_{m-1}g^{r_{m}})^{r_{m-1}}\cdots)^{r_{1}})^{r_{0}},g_{0}(g_{1}(\cdots(g_{m})^{r_{m-1}}\cdots)^{r_{1}})^{r_{0}})<\frac{1}{3^{m}}$.
Select $g_{m+1}\in U$ such that $\phi(g_{m+1})$ has order greater than
$((m+1)r_{0}\cdots r_{m})!\cdot r_{0}\cdot r_{1}\cdots\cdot r_{m}.$
Let $r_{m+1}$ be the least common multiple of the finite nonempty set $\{\operatorname{ex}(H_{j})|j\in\textbf{supp}(\phi(g_{m+1}))\}$. Now our sequences are defined.
For $m\leq k\in\omega$ define the element
$g_{m,k}=g_{m}(\cdots(g_{k-1}(g_{k})^{r_{k-1}})^{r_{k-2}}\cdots)^{r_{m}}.$
For $m\leq s<t\in\omega$ we have by construction
$\begin{array}[]{ll}d(g_{m,s},g_{m,t})&\leq d(g_{m,s},g_{m,s+1})+d(g_{m,s+1},g_{m,s+2})+\cdots+d(g_{m,t-1},g_{m,t})\vspace*{2mm}\\ &<\frac{1}{3^{s+1-m}}+\frac{1}{3^{s+2-m}}+\cdots+\frac{1}{3^{t-m}}\end{array}$
In particular the sequence $g_{m,m},g_{m,m+1},g_{m,m+1},\ldots$ is Cauchy and converges to, say, $g_{m,\infty}$. By continuity we see that for each $m\in\omega$ the equation
$g_{m,\infty}=g_{m}g_{m+1,\infty}^{r_{m}}$
holds.
We know that $\phi(g_{0,\infty})$ has finite order and so let $N\in\omega\setminus\{0\}$ be such that $(\phi(g_{0,\infty}))^{N}=1_{H}$. As $(\phi(g_{N}))^{r_{0}\cdots r_{N-1}}$ has order greater than $(N\cdot r_{0}\cdots r_{N-1})!$, select $j_{0}\in\textbf{supp}(g_{N})$ such that $\pi_{j_{0}}(\phi(g_{N}))^{r_{0}\cdots r_{N-1}}$ has order greater than $N\cdot r_{0}\cdots r_{N-1}$, where $\pi_{j_{0}}:\prod_{j\in J}H_{j}\rightarrow H_{j_{0}}$ is projection. Clearly $\pi_{j_{0}}\circ\phi(g_{N}(g_{N+1,\infty})^{r_{N}}))=\pi_{j_{0}}\circ\phi(g_{N})$ since $\operatorname{ex}(H_{j_{0}})$ divides $r_{N}$. If $j_{0}\in\bigcup_{\ell=0}^{N-1}\textbf{supp}(g_{\ell})$ then $\operatorname{ex}(H_{j_{0}})$ divides $r_{k}$ for some $10\leq k\leq N-$ 1, so in particular $\pi_{j_{0}}\circ\phi(g_{N}^{r_{0}\cdots r_{N-1}})=1_{H_{j_{0}}}$, which is a contradiction. Therefore
$\pi_{j_{0}}\circ\phi(g_{0,\infty})=\pi_{j_{0}}\circ\phi(g_{0}(g_{1}(\cdots g_{N-1}(g_{N}g_{N+1,\infty}^{r_{N}})^{r_{N-1}}\cdots)^{r_{1}})^{r_{0}})=\pi_{j_{0}}\circ\phi(g_{N}^{r_{0}\cdots r_{N-1}})$
has order greater than $N\cdot r_{0}\cdots r_{N-1}>N$, while $\phi(g_{0,\infty})$ had order $N$, contradiction.
In situation (2) we construct
  1. (i)
    a sequence $(g_{m})_{m\in\omega}$ of elements in $G$;
  2. (ii)
    a sequence $(r_{m})_{m\in\omega}$ of positive natural numbers; and
  3. (iii)
    a sequence $(V_{m})_{m\in\omega}$ of open neighborhoods of $1_{G}$.
Pick $g_{0}\in G$ such that $\phi(g_{0})$ has order greater than $1$, take $r_{0}$ to be the least common multiple of $\{\operatorname{ex}(H_{j})|j\in\textbf{supp}(\phi(g_{0}))\}$, and let $V_{0}$ be an open neighborhood of $1_{G}$ such that $\overline{V_{0}}$ is countably compact. Assume we have made the selections for all subscripts less than or equal to $m$. Select $g_{m+1}\in V_{m}$ such that $\phi(g_{m+1})$ has order greater than $((m+1)r_{0}\cdots r_{m})!\cdot r_{0}\cdot r_{1}\cdots\cdot r_{m}$. Let $r_{m+1}$ be the least common multiple of $\{\operatorname{ex}(H_{j})|j\in\textbf{supp}(\phi(g_{m+1}))\}$. Let $V_{m+1}$ be a neighborhood of $1_{G}$ such that
$g_{m+1}V_{m+1}^{r_{m+1}}\subseteq V_{m}.$
Note that $g_{m+1}(\overline{V_{m+1}})^{r_{m+1}}\subseteq\overline{V_{m}}$ for each $m\in\omega$. Letting
$Y_{m}=g_{0}(g_{1}(g_{2}(\cdots g_{m}(\overline{V_{m}})^{r_{m}}\cdots)^{r_{2}})^{r_{1}})^{r_{0}}$
for $m\in\omega$, we have $Y_{m}\supseteq Y_{m+1}$ and $Y_{m}$ is countably compact. Select $g_{\infty}\in\bigcap_{m\in\omega}Y_{m}$ and let $N$ be the order of $\phi(g_{\infty})$. Select $\underline{g}\in\overline{V_{N}}$ for which
$g_{\infty}=g_{0}(g_{1}(g_{2}(\cdots g_{N}(\underline{g})^{r_{N}}\cdots)^{r_{2}})^{r_{1}})^{r_{0}}.$
Select $j_{0}$ and derive the same contradiction as in situation (1). ∎
Corollary 22.
Suppose $\{G_{i}\}_{i\in\omega}$ is a collection of groups, $\{H_{j}\}_{j\in J}$ is a collection of bounded torsion groups, and $\mathcal{U}$ is a nonprincipal ultrafilter on $\omega$. Then the image of any abstract homomorphism
$\phi:(\prod_{\omega}G_{i})/\mathcal{U}\rightarrow\textbf{Bdd}\{H_{j}\}_{j\in J}$
is a bounded torsion group.
Proof.
Assume the hypotheses. By considering each group $G_{i}$ as a discrete topological group, it is well-known that the topological group $\prod_{i\in\omega}G_{i}$ is completely metrizable, and the subgroups of form $\{1_{G_{0}}\}\times\cdots\times\{1_{G_{M}}\}\times\prod_{i>M}G_{i}$ give a basis of neighborhoods for identity. Applying Theorem 21 we have some $M,N\in\omega$ for which $\phi\circ\rho(\{1_{G_{0}}\}\times\cdots\times\{1_{G_{M}}\}\times\prod_{i>M}G_{i})$ consists of elements of order dividing $N$, where $\rho:\prod_{\omega}G_{i}\rightarrow(\prod_{\omega}G_{i})/\mathcal{U}$ is the natural map. Since the restriction $\rho\upharpoonright\{1_{G_{0}}\}\times\cdots\times\{1_{G_{M}}\}\times\prod_{i>M}G_{i}$ is surjective onto the ultraproduct, we are done. ∎
For convenience, in the remainder of this section we’ll say a group which is a homomorphic image of a nonprincipal ultraproduct over $\omega$ is a ui-group (“ui” for ultraproduct image).
Corollary 23.
If $\{H_{j}\}_{j\in J}$ is a collection of bounded torsion groups for which $\{\operatorname{ex}(H_{j})\}_{j\in J}$ is unbounded in $\omega$, then the direct sum $\bigoplus_{j\in J}H_{j}$ is not a ui-group.
Part of the attraction of Corollary 23 is that it provides many examples of non-ui-groups which are (unbounded) torsion. One can apply Corollary 23 to the collection $\{S_{j}\}_{j\in\omega}$ where $S_{j}$ is the symmetric group on $j$ elements, or to the collection $\{\mathfrak{b}(2,j)\}_{j\in\omega\setminus\{0,1\}}$ where $\mathfrak{b}(2,j)$ is the free Burnside group of rank $2$ and exponent $j$.
There exist ui-groups which are unbounded torsion. For example, a quasicyclic group $\mathbb{Z}(p^{\infty})$ is cotorsion and therefore a ui-group. We do not know the answers to the following.
Questions.
  1. (A)
    If a ui-group is unbounded torsion then does it include a nontrivial divisible subgroup?
  2. (B)
    Is there a bounded torsion group which is not a ui-group?
A classical open problem asks whether there exists a compact unbounded torsion group $C$ [12, Qu. 17.93]. Such a $C$ would be profinite [11, Theorem 28.20] and therefore any divisible subgroup would be trivial. Moreover for any nonprincipal ultrafilter $\mathcal{U}$ on $\omega$ it is easy to construct a homomorphism from the ultrapower $(\prod_{\omega}C)/\mathcal{U}$ onto $C$. Thus $C$ would give a negative answer to Question (A).
We give a few more examples of non-ui-groups. Recall that a group $H$ is cm-slender if every abstract group homomorphism from a completely metrizable group to $H$ has open kernel [6]. A cm-slender group is necessarily torsion-free, and examples include Baumslag-Solitar groups and nontrivial word-hyperbolic groups. Arguing as in Corollary 22 it is clear that a group having nontrivial homomorphic image in a cm-slender group is not a ui-group. Reasoning similarly, an infinite subgroup of the mapping class group of a connected compact surface is not a ui-group [4, Theorem 9.1], nor is a group which is nontrivially a free product of groups [13, Theorem 1.5].
Remark 24.
By making easy changes to the proof of Theorem 21 one can show an analogous statement where $G$ is instead the fundamental group $\pi_{1}(E)$ of the infinite earring $E$ [5].
4. A common element below two ultrafilters
In this brief section we remind the reader of the Rudin-Keisler ordering (see e.g. [9, Ch. 11]) and point out why Question 5 has a consistent affirmative answer.
Definition 25.
For ultrafilters $\mathcal{U}$ and $\mathcal{U}^{\prime}$ on $\omega$ we write $\mathcal{U}\leq_{RK}\mathcal{U}^{\prime}$ if there exists a function $f:\omega\rightarrow\omega$ such that $f(\mathcal{U}^{\prime}):=\{X\subseteq\omega\mid(\exists Y\in\mathcal{U}^{\prime})X\supseteq f(Y)\}=\mathcal{U}$.
Proposition 26.
Suppose that $\mathcal{U}\leq_{RK}\mathcal{U}^{\prime}$. Then every homomorphic image of an ultraproduct over $\mathcal{U}^{\prime}$ is also a homomorphic image of an ultraproduct over $\mathcal{U}$.
Proof.
Assume the hypotheses. Let $(G_{i})_{i\in\omega}$ be a sequence of groups. Take $f:\omega\rightarrow\omega$ such that $f(\mathcal{U}^{\prime})=\mathcal{U}$. For each $j\in\omega$ in the image of $f$ select a cardinal $\kappa_{j}$ large enough that there is a homomorphic surjection $\phi_{j}:F_{\kappa_{j}}\rightarrow\prod_{i\in f^{-1}(\{j\})}G_{i}$ from the free group $F_{\kappa_{j}}$ of rank $\kappa_{j}$. For $j\in\omega$ which is not in the image of $f$ we let $\kappa_{j}=0$. Now we obtain a composition $\Psi=\rho\circ\Phi$ of homomorphisms
${\prod_{j\in\omega}F_{\kappa_{j}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${\prod_{i\in\omega}G_{i}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho}$${(\prod_{i\in\omega}G_{i})/\mathcal{U}^{\prime}}$
which are evidently surjective ($\Phi$ is defined componentwise by the $\phi_{j}$). Suppose that $(W_{j})_{j\in\omega}\in\prod_{\omega}F_{\kappa_{j}}$ is such that $W_{j}$ is identity for almost every $j$, say for all elements $j$ in $X\in\mathcal{U}$. Then for all $j\in X$ we have $\phi_{j}(W_{j})$ is identity, and as $f^{-1}(X)\in\mathcal{U}^{\prime}$, we have $(W_{j})_{j\in\omega}$ in the kernel of $\Psi$. So, $\Psi$ descends to a homomorphic surjection $\overline{\Psi}:(\prod_{j\in\omega}F_{\kappa_{j}})/\mathcal{U}\rightarrow(\prod_{i\in\omega}G_{i})/\mathcal{U}^{\prime}$. Thus every homomorphic image of $(\prod_{i\in\omega}G_{i})/\mathcal{U}^{\prime}$ is also a homomorphic image of $(\prod_{j\in\omega}F_{\kappa_{j}})/\mathcal{U}$. ∎
The assertion Near coherence of filters (NCF) states that for any two nonprincipal ultrafilters $\mathcal{U}$ and $\mathcal{U}^{\prime}$ there exists a finite-to-one function $f:\omega\rightarrow\omega$ such that $f(\mathcal{U})=f(\mathcal{U}^{\prime})$ [2, §5]. In particular, for any two nonprincipal ultrafilters $\mathcal{U}$ and $\mathcal{U}^{\prime}$, NCF gives a nonprincipal ultrafilter $\mathcal{U}^{\prime\prime}$ which is $\leq_{RK}$ below both $\mathcal{U}$ and $\mathcal{U}^{\prime}$. By Proposition 26 the homomorphic images of ultraproducts over $\mathcal{U}$ or $\mathcal{U}^{\prime}$ are also homomorphic images of ultraproducts over $\mathcal{U}^{\prime\prime}$. Since NCF is consistent with ZFC [3], Question 5 has a consistent positive answer.
Acknowledgement
The author thanks the referees for pointing out clarifications and typos in a very timely manner.